# Multitask-based Evaluation of Open-Source LLM on Software Vulnerability

### v2

@@ 3.6 Implementation
We develop the generation pipeline in Python, utilizing PyTorch [69] implementations of DeepSeek Coder, CodeLlama, StarCoder, WizardCoder, Mistral, and Phi-2. We use the Huggingface [70] to load the model weights and generate outputs. We also adhere to the best-practice guide [71] for each prompt. For the fine-tuning setting, we select the models with 7B parameters or less, and for the few-shot setting, we use models with fewer than 34B parameters. To directly compare the fine-tuning setting with the few-shot setting, we employ models with the same parameter in both settings (i.e., DeepSeek Coder 6.7B, CodeLlama 7B, StarCoder 7B, WizardCoder 7B, Mistral 7B, and Phi-2 2.7B). The constraint on the number of parameters is imposed by our computing resources. Table III summarizes the characteristics of the studied LLMs. Furthermore, considering the limitation of LLM’s conversation windows, we manually select three examples for the few-shot setting from the training set. Regarding ReVeal, IVDetect, Devign, LineVul, and SVulD, we utilize their publicly available source code and perform fine-tuning with the default parameters provided in their original code. Considering Devign’s code is not publicly available, we make every effort to replicate its functionality and achieve similar results on the original paper’s dataset. All these models are implemented using the PyTorch [69] framework. The evaluation is conducted on a 16-core workstation equipped with an Intel(R) Xeon(R) Gold 6226R CPU @ 2.90Ghz, 192GB RAM, and 10 × NVIDIA RTX 3090 GPU, running Ubuntu 20.04.1 LTS.
@@ 4 Experimental results
4.1 RQ-1: Evaluating Vulnerability Detection of LLMs
We consider the five SOTA baselines: Devign [36], ReVeal [45], IVDetect [52], LineVul [37], and SVulD [46]. These approaches can be divided into two groups: graph-based (i.e., Devign, ReVeal and IVDetect) and transformer-based (i.e., LineVul and SVulD). Besides, in order to comprehensively compare the performance among baselines and LLMs, we consider four widely used performance measures (i.e., Precision, Recall, F1-score, and Accuracy) and conduct experiments on the popular dataset. Since graph-based approaches need to obtain the structure information (e.g., control flow graph (CFG), data flow graph (DFG)) of the studied functions, we adopt the same toolkit with Joern to transform functions. The functions are dropped out directly if they cannot be transformed by Joern successfully. Finally, the filtered dataset (shown in Table II) is used for evaluation. We follow the same strategy to build the training data, validating data, and testing data from the original dataset with previous work does [37, 54]. Specifically, 80% of functions are treated as training data, 10% of functions are treated as validation data, and the left 10% of functions are treated as testing data. We also keep the distribution as same as the original ones in training, validating, and testing data. Apart from presenting the overall performance comparison, we also give the detailed performance of LLMs on the Top-10 CWE types for a better analysis.
Results. [A] LLMs vs. SOTA approaches. Table V shows the overall performance measures between LLMs and five baselines and the best performances are highlighted in bold. According to the results in Table V, we can obtain the following observations:
(1) Fine-tuned LLMs have poor performance compared with transformer-based approaches when considering F1-score, Precision, and Accuracy. In particular, SVulD obtains 0.336, 0.282, and 0.915 in terms of F1-score, Precision, and Accuracy, which surpass the fine-tuned LLMs by 23.5%-242.9%, 70.9%-432.1%, and 16.9%-162.2% in terms of F1-score, Precision, and Accuracy, respectively. LineVul also outperforms all LLMs in precision and accuracy, with F1-score equal to the best LLM.
(2) The performance of fine-tuned LLMs is comparable to graph-based approaches. As for recall, LLMs under the fine-tuning setting generally outperform previous SOTA approaches, except for StarCoder (0.646), which falls slightly behind the top-performing approach, Devign (0.660). As for F1-score, LLMs achieve a range of 0.214-0.272, whereas graph-based approaches achieve a range of 0.200-0.232. As for Precision, LLMs score between 0.123 and 0.159, while graph-based methods range from 0.118 to 0.172. Additionally, in terms of Accuracy, LLMs range from 0.685 to 0.771, whereas graph-based methods range from 0.726 to 0.815.
(3) LLMs under few-shot setting have poor performance compared with existing approaches. LLMs ranging from 2.7B to 34B parameters perform less favorably than existing approaches in terms of F1-score and Precision. However, as for Accuracy, SVulD (transformer-based) obtains the best performance (0.915) and DeepSeek-Coder 6.7B ranks third (0.823), which is better than the three graph-based approaches.
Finding-1. LLMs can detect software vulnerabilities, but fine-tuned LLMs perform weaker than transformer-based methods, yet comparably to graph-based methods. Moreover, LLMs in the few-shot setting show lower performance than existing methods.
[B] Fine-Tuning vs. Few-Shot. The experimental results are presented in Table  V. Based on these experimental findings, we can draw the following observations: (1) LLMs fine-tuned for vulnerability detection demonstrate superior performance on the task compared to LLMs (except Mistral) in the few-shot setting. The F1-score and Precision have doubled, while the Recall has also shown improvement. (2) Mistral experiences a significant drop in performance after fine-tuning, except for Recall. This could be due to the absence of high-quality security vulnerability-related data in its pre-training. (3) LLMs with more parameters typically exhibit better performance. For example, CodeLlama 34B improves upon CodeLlama 7B by 19.4%, 34.5%, and 37.0% in terms of F1-score, Precision, and Accuracy, respectively. However, different LLMs may exhibit performance variations due to differences in model design and the quality of pre-training data. (4) Phi-2 achieves performance approximating that of other LLMs with 7 billion parameters, even with a parameter size of 2.7 billion. This may be attributed to the higher quality of its pre-training data.
Finding-2. After fine-tuning, the detection capability of LLMs has improved, except for Mistral. Larger models usually perform better, but performance can also be influenced by model design and pre-training data.
[C] The comparisons of Top-10 CWE types between LLMs. Table VI shows the detailed comparisons of Top-10 CWE types between fine-tuned LLMs. In this table, we highlight the best performance for each performance metric in bold. According to the results, we can achieve the following observations:
(1) In most cases, WizardCoder obtains better performance than LLMs by considering all performance metrics. Mistral performs the worst on many CWE types, even weaker than other LLMs by several times, which means Mistral has almost no ability to detect vulnerabilities. Other LLMs have certain advantages in different CWE types, complementing each other.
(2) Considering the performance of all metrics, WizardCoder achieves the best performances on CWE-119 (“Improper Restriction of Operations within the Bounds of a Memory Buffer”), CWE-125 (“Out-of-bounds Read”), CWE-362 (“Concurrent Execution using Shared Resource with Improper Synchronization (’Race Condition’)”), and CWE-476 (“NULL Pointer Dereference”), which indicates WizardCoder is exceptionally skilled at detecting and mitigating vulnerabilities related to memory handling and synchronization issues.
Finding-3. WizardCoder has the best vulnerability detection capability, while Mistral is the worst, with the remaining LLMs complementing each other.
@@ 4.1 RQ-1: Evaluating Vulnerability Detection of LLMs
TABLE V: The comparison between LLMs and five baselines on software vulnerability detection
@@
DeepSeek-Coder 6.7B 0.214 0.830 0.123 0.685
CodeLlama 7B 0.265 0.797 0.159 0.771
StarCoder 7B 0.224 0.646 0.136 0.769
WizardCoder 7B 0.272 0.781 0.165 0.783
Mistral 0.098 0.686 0.053 0.349
Phi-2 0.239 0.754 0.142 0.751
@@
TABLE VI: The software vulnerability detection comparison on Top-10 CWEs among fine-tuned LLMs
@@
CWE-119 1549 128 0.742 0.680 0.680 0.758 0.633 0.719 0.701 0.750 0.751 0.757 0.349 0.747
CWE-20 1082 80 0.828 0.859 0.734 0.813 0.797 0.766 0.723 0.753 0.756 0.755 0.369 0.731
CWE-264 800 64 0.863 0.788 0.700 0.888 0.738 0.800 0.753 0.803 0.786 0.819 0.356 0.784
CWE-399 697 35 0.889 0.815 0.852 0.852 0.741 0.963 0.785 0.819 0.829 0.845 0.340 0.798
CWE-125 582 29 0.828 0.690 0.655 0.828 0.724 0.828 0.704 0.751 0.771 0.775 0.345 0.771
CWE-200 573 27 0.743 0.686 0.629 0.714 0.657 0.629 0.731 0.782 0.763 0.810 0.328 0.763
CWE-189 442 21 0.688 0.750 0.625 0.875 0.625 0.438 0.742 0.817 0.776 0.833 0.367 0.753
CWE-362 413 16 0.800 0.600 0.600 0.600 0.600 0.400 0.714 0.792 0.758 0.782 0.349 0.753
CWE-416 406 12 0.667 0.500 0.333 0.583 0.500 0.667 0.719 0.778 0.786 0.781 0.323 0.724
CWE-476 367 11 0.857 0.714 0.429 1.000 0.429 0.714 0.665 0.728 0.684 0.738 0.381 0.695
@@ 4.2 RQ-2: Evaluating Vulnerability Assessment of LLMs
TABLE VII: The comparison of LLMs on software vulnerability assessment
@@
DeepSeek-Coder 6.7B 0.573 0.547 0.648 0.691
CodeLlama 7B 0.621 0.592 0.684 0.703
StarCoder 7B 0.563 0.539 0.614 0.667
WizardCoder 7B 0.558 0.537 0.619 0.686
Mistral 0.318 0.343 0.329 0.541
Phi-2 0.495 0.481 0.599 0.663
@@
DeepSeek-Coder 6.7B 0.244 0.375 0.358 0.253
DeepSeek-Coder 33B 0.302 0.342 0.338 0.352
CodeLlama 7B 0.303 0.331 0.319 0.367
CodeLlama 34B 0.274 0.367 0.321 0.301
StarCoder 7B 0.314 0.363 0.358 0.450
StarCoder 15.5B 0.309 0.378 0.340 0.356
WizardCoder 7B 0.238 0.357 0.334 0.265
WizardCoder 34B 0.334 0.388 0.368 0.390
Mistral 0.201 0.349 0.320 0.202
Phi-2 0.266 0.346 0.335 0.312
TABLE VIII: A vulnerable code for CodeLlama to assess with different prompts
@@
Results. Table VII shows the detailed results of LLMs on vulnerable assessment. Based on these experimental results, we can observe a significant improvement in the vulnerability assessment capability of LLMs after fine-tuning. Specifically, the accuracy has increased from 0.202-0.471, reaching a range of 0.541-0.703, while precision has improved from 0.319-0.358, now ranging from 0.329-0.684. This underscores the necessity of using fine-tuning in downstream tasks. It’s worth noting that CodeLlama, after fine-tuning, achieves the best performance across all metrics. If researchers need to perform tasks such as vulnerability assessment with LLM, we recommend choosing CodeLlama as the preferred option. We also find that Mistral exhibits a relatively smaller improvement after fine-tuning, which aligns with our expectations, as it is a general LLM. In a few-shot setting, StarCoder stands out by achieving performance comparable to 34B LLMs with a significantly smaller parameter size (i.e., 7B). In this context, it achieves F1-score and Accuracy metrics of 0.314 and 0.450, respectively.
Finding-4: Larger model parameter counts did not enhance vulnerability assessment performance with LLMs, prioritize smaller parameter models for better cost-performance balance.
@@
Case Study. To illustrate the effectiveness of key important information, we present an instance of a vulnerability (CWE-119) in Big-Vul that is exclusively assess by CodeLlama, as depicted in Table VIII. This example is a vulnerability in the Linux project, categorized under CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer Vulnerability). In an initial assessment without critical information, CodeLlama did not fully grasp the severity of this vulnerability and labeled it as “Medium”. However, with the provision of crucial details, CodeLlama can more accurately evaluate the risk level of this vulnerability. The CVE description for this vulnerability highlights multiple buffer overflows in the net/wireless/nl80211.c file of the Linux kernel prior to version 2.6.39.2. These vulnerabilities allow local users to gain elevated privileges by leveraging the CAP NET ADMIN capability during scan operations with an excessively long SSID value. In this scenario, the lack of proper validation of the SSID length leads to buffer overflows, enabling attackers to exploit the vulnerability, escalate privileges, and execute malicious code. The commit message described that this bug has existed since version 2.6.29-rc4 of the Linux kernel. Given this information, CodeLlama reassesses the risk level of this vulnerability as “High”. This is because it allows attackers to escalate privileges and execute malicious code, and it has persisted for a considerable period of time. It is crucial to address and patch this vulnerability promptly by updating the operating system or kernel to ensure security.
@@
To compare the vulnerability assessment capabilities of LLMs after providing key information, we have created a performance comparison bar chart, as shown in Fig.4. LLMs have limited capacity for assessing vulnerability severity based solely on source code. However, when provided with key important information, most LLMs (i.e., DeepSeek-Coder, CodeLlama, WizardCoder, and Mistral) exhibit significantly improved vulnerability assessment capabilities, particularly in terms of the Accuracy metric. The Accuracy has increased from the range of 0.2-0.37 to the range of 0.21-0.43. StarCoder and Phi-2 are showing a declining trend, and we believe this may be attributed to the addition of key information, resulting in an increase in the number of input tokens. These LLMs may not excel in handling excessively long text sequences. In contrast, DeepSeek-Coder exhibits a significant improvement, possibly due to its proficiency in handling long sequential text.
Figure 4: The impact of key important information on LLM Vulnerability Assessment
@@
Finding-5: LLMs have the limited capacity for assessment of vulnerability severity based on source code only, but can be extremely improved if provided with more context information in most cases.

### v3

@@ 3.7 Implementation
We develop the generation pipeline in Python, utilizing PyTorch [75] implementations of DeepSeek Coder, CodeLlama, StarCoder, WizardCoder, Mistral, and Phi-2. We use the Huggingface [76] to load the model weights and generate outputs. We also adhere to the best-practice guide [77] for each prompt. For the fine-tuning setting, we select the models with 7B parameters or less, and for the few-shot setting, we use models with fewer than 34B parameters. To directly compare the fine-tuning setting with the few-shot setting, we employ models with the same parameter in both settings (i.e., DeepSeek-Coder 6.7B, CodeLlama 7B, StarCoder 7B, WizardCoder 7B, Mistral 7B, and Phi-2 2.7B). The constraint on the number of parameters is imposed by our computing resources. Table III summarizes the characteristics of the studied LLMs. Furthermore, considering the limitation of LLM’s conversation windows, we manually select three examples for the few-shot setting from the training data. Regarding baselines (i.e., pre-trained LMs, Reveal, IVDetect, Devign, LineVul, and SVulD), we utilize their publicly available source code and perform fine-tuning with the default parameters provided in their original code. Considering Devign’s code is not publicly available, we make every effort to replicate its functionality and achieve similar results on the original paper’s dataset. All these models are implemented using the PyTorch [75] framework. The evaluation is conducted on a 16-core workstation equipped with an Intel(R) Xeon(R) Gold 6226R CPU @ 2.90Ghz, 192GB RAM, and 10 × NVIDIA RTX 3090 GPU, running Ubuntu 20.04.1 LTS.
@@ 4 Experimental results
TABLE VI: The comparison between LLMs and eleven baselines on software vulnerability detection (RQ1)
@@
CodeBERT 0.270 0.608 0.173 0.830
GraphCodeBERT 0.246 0.721 0.148 0.771
UniXcoder 0.256 0.787 0.153 0.764
PLBART 0.255 0.692 0.157 0.791
CodeT5 0.237 0.759 0.141 0.748
CodeT5+ 0.218 0.508 0.139 0.812
@@
DeepSeek-Coder 6.7B 0.270 0.627 0.172 0.824
CodeLlama 7B 0.259 0.806 0.154 0.761
StarCoder 7B 0.220 0.607 0.135 0.778
WizardCoder 7B 0.214 0.365 0.151 0.861
Mistral 0.220 0.607 0.135 0.778
Phi-2 0.241 0.557 0.154 0.818
@@
TABLE VII: The software vulnerability detection comparison on Top-10 CWEs among fine-tuned LLMs (RQ1)
@@
CWE-119 1549 128 0.570 0.719 0.625 0.359 0.453 0.492 0.801 0.735 0.777 0.839 0.785 0.792
CWE-20 1082 80 0.609 0.844 0.609 0.313 0.297 0.578 0.804 0.735 0.758 0.866 0.793 0.814
CWE-264 800 64 0.763 0.900 0.613 0.400 0.525 0.700 0.839 0.795 0.759 0.850 0.810 0.846
CWE-399 697 35 0.815 0.926 0.630 0.481 0.556 0.704 0.885 0.821 0.815 0.901 0.854 0.877
CWE-125 582 29 0.586 0.931 0.621 0.414 0.517 0.517 0.808 0.746 0.771 0.830 0.763 0.765
CWE-200 573 27 0.514 0.743 0.600 0.286 0.343 0.457 0.829 0.743 0.770 0.841 0.784 0.812
CWE-189 442 21 0.625 0.813 0.813 0.563 0.500 0.750 0.853 0.776 0.828 0.891 0.833 0.869
CWE-362 413 16 0.200 0.800 0.600 0.200 0.200 0.200 0.847 0.794 0.821 0.908 0.821 0.855
CWE-416 406 12 0.667 0.750 0.667 0.333 0.500 0.500 0.835 0.796 0.773 0.884 0.828 0.820
CWE-476 367 11 0.571 1.000 0.429 0.286 0.286 0.143 0.782 0.687 0.706 0.820 0.719 0.725
4.1 RQ-1: Evaluating Vulnerability Detection of LLMs
In addition to pre-trained LMs, we also consider the following five SOTA baselines: Devign [38], Reveal [47], IVDetect [56], LineVul [39], and SVulD [48]. These baselines can be divided into two groups: graph-based (i.e., Devign, Reveal, and IVDetect) and transformer-based (i.e., pre-trained LMs, LineVul, and SVulD). Besides, in order to comprehensively compare the performance among baselines and LLMs, we consider four widely used performance measures (i.e., Precision, Recall, F1-score, and Accuracy) and conduct experiments on the popular dataset. Since graph-based approaches need to obtain the structure information (e.g., control flow graph (CFG), data flow graph (DFG)) of the studied functions, we adopt the same toolkit with Joern to transform functions. The functions are dropped out directly if they cannot be transformed by Joern successfully. Finally, the filtered dataset (shown in Table II) is used for evaluation. We follow the same strategy to build the training data, validation data, and testing data from the original dataset with previous work does [39, 58]. Specifically, 80% of functions are treated as training data, 10% of functions are treated as validation data, and the left 10% of functions are treated as testing data. We also keep the distribution as same as the original ones in training, validation, and testing data. We undersample the non-vulnerable functions to produce approximately balanced training data at the function level, while the validation and testing data remain in the original imbalanced ratio. Apart from presenting the overall performance comparison, we also give the detailed performance of LLMs on the Top-10 CWE types for a better analysis.
Results. [A] LLMs vs. Baselines. Table VI shows the overall performance measures between LLMs and eleven baselines and the best performances are highlighted in bold. According to the results in Table VI, we can obtain the following observations:
(1) Fine-tuned LLMs have poor performance compared with transformer-based approaches when considering F1-score, Precision, and Accuracy. In particular, SVulD obtains 0.336, 0.282, and 0.915 in terms of F1-score, Precision, and Accuracy, which surpass the fine-tuned LLMs by 24.4% to 57.0%, 64.0% to 108.9%, and 6.3% to 20.2% in terms of F1-score, Precision, and Accuracy, respectively. Notably, the F1-score performance of LineVul is significantly lower (0.272) than that reported in the original paper (0.910). We further analyze this discrepancy in Section 5.1.
(2) The performance of fine-tuned LLMs is comparable to graph-based approaches. For example, in terms of F1-score, fine-tuned LLMs achieve a range of 0.214 to 0.270. In comparison, graph-based approaches achieve a range of 0.200 to 0.232.
(3) LLMs under few-shot setting have poor performance compared with baselines. LLMs ranging from 2.7B to 34B parameters perform less favorably than baselines in terms of F1-score and Precision. However, as for Accuracy, SVulD (transformer-based) obtains the best performance (0.915) and DeepSeek-Coder 6.7B under few-shot setting achieves a performance of 0.823, which is better than the three graph-based approaches.
Finding-1. LLMs can detect vulnerabilities, but fine-tuned LLMs perform weaker than transformer-based approaches. Considering the computational resources and time costs of deploying LLMs, transformer-based approaches for vulnerability detection are a more efficient choice.
[B] Fine-Tuning vs. Few-Shot. The experimental results are presented in Table VI. Based on these experimental findings, we can draw the following observations: (1) LLMs fine-tuned for vulnerability detection demonstrate superior performance on the task compared to LLMs in the few-shot setting. The average F1-score and average Precision have doubled, while the average Recall has also shown improvement. (2) LLMs with more parameters typically exhibit better performance. For example, CodeLlama 34B improves upon CodeLlama 7B by 19.4%, 34.5%, and 37.0% in terms of F1-score, Precision, and Accuracy, respectively. However, different LLMs may exhibit performance variations due to differences in model design and the quality of pre-training data. (3) Phi-2 achieves performance approximating that of other LLMs with 7 billion parameters, even with a parameter size of 2.7 billion. This may be attributed to the higher quality of its pre-training data.
Finding-2. After fine-tuning, the detection capability of LLMs has improved. Larger models usually perform better, but performance can also be influenced by model design and pre-training data. Therefore, fine-tuning the LLM on domain-specific data before using it as a vulnerability detector is necessary.
[C] The comparisons of Top-10 CWE types between LLMs. Table VII shows the detailed comparisons of Top-10 CWE types between fine-tuned LLMs. In this table, we highlight the best performance for each performance metric in bold. According to the results, we can achieve the following observations: (1) In most cases, CodeLlama obtains better performance than other LLMs in terms of F1-score, Precision, and Recall. Different LLMs have certain advantages in different CWE types, complementing each other. (2) Considering the performance of F1-score, Precision, and Recall, CodeLlama achieves the best performances on CWE-125 (“Out-of-bounds Read”), CWE-362 (“Concurrent Execution using Shared Resource with Improper Synchronization (’Race Condition’)”), and CWE-476 (“NULL Pointer Dereference”), which indicates CodeLlama is exceptionally skilled at detecting and mitigating vulnerabilities related to memory handling and synchronization issues.
Finding-3. In general, different LLMs complementing each other, while CodeLlama obtains better performance in terms of F1-score, Precision, and Recall.
@@ 4.2 RQ-2: Evaluating Vulnerability Assessment of LLMs
Results. Table VIII shows the detailed results of LLMs and six baselines on vulnerability assessment. Based on these experimental results, we can observe a significant improvement in the vulnerability assessment capability of LLMs after fine-tuning. Specifically, the accuracy has increased from 0.282 to 0.424, reaching a range of 0.759 to 0.860, while precision has improved from 0.296 to 0.355, now ranging from 0.512 to 0.854. This underscores the necessity of fine-tuning in vulnerability assessment task. Overall, fine-tuned code-related LLMs outperform pre-trained LMs in vulnerability assessment. It is worth noting that DeepSeek-Coder, after fine-tuning, achieves the best performance compared to other LLMs and pre-trained LMs. If researchers need to perform tasks such as vulnerability assessment with LLM, fine-tuning DeepSeek-Coder is a more efficient choice. We also find that Mistral exhibits a relatively smaller improvement after fine-tuning, which aligns with our expectations, as it is a general LLM.
TABLE VIII: The comparison between LLMs and six baselines on software vulnerability assessment (RQ2)
@@
CodeBERT 0.753 0.730 0.788 0.828
GraphCodeBERT 0.701 0.666 0.772 0.802
UniXcoder 0.745 0.761 0.734 0.817
PLBART 0.735 0.741 0.731 0.789
CodeT5 0.743 0.750 0.741 0.817
CodeT5+ 0.706 0.677 0.755 0.789
@@
DeepSeek-Coder 6.7B 0.814 0.785 0.854 0.860
CodeLlama 7B 0.768 0.749 0.794 0.827
StarCoder 7B 0.671 0.677 0.666 0.764
WizardCoder 7B 0.793 0.778 0.813 0.842
Mistral 0.525 0.539 0.512 0.759
Phi-2 0.747 0.732 0.767 0.802
@@
DeepSeek-Coder 6.7B 0.229 0.339 0.310 0.262
DeepSeek-Coder 33B 0.290 0.323 0.336 0.335
CodeLlama 7B 0.310 0.331 0.334 0.373
CodeLlama 34B 0.265 0.323 0.327 0.294
StarCoder 7B 0.265 0.342 0.333 0.330
StarCoder 15.5B 0.285 0.315 0.329 0.326
WizardCoder 7B 0.244 0.351 0.336 0.250
WizardCoder 34B 0.306 0.330 0.325 0.379
Mistral 0.283 0.308 0.296 0.424
Phi-2 0.269 0.359 0.355 0.282
TABLE IX: A vulnerable code for CodeLlama to assess with different prompts (RQ2)
@@
Finding-4. Overall, fine-tuned code-related LLMs outperform pre-trained LMs in vulnerability assessment. When resources permit, fine-tuning DeepSeek-Coder 6.7B for vulnerability assessment is optimal, as it outperforms the pre-trained LMs across four metrics.
@@
Case Study. To illustrate the effectiveness of key important information, we present an instance of a vulnerability (CWE-119) in Big-Vul that is exclusively assess by CodeLlama, as depicted in Table IX. This example is a vulnerability in the Linux project, categorized under CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer Vulnerability). In an initial assessment without critical information, CodeLlama did not fully grasp the severity of this vulnerability and labeled it as “Medium”. However, with the provision of crucial details, CodeLlama can more accurately evaluate the risk level of this vulnerability. The CVE description for this vulnerability highlights multiple buffer overflows in the net/wireless/nl80211.c file of the Linux kernel prior to version 2.6.39.2. These vulnerabilities allow local users to gain elevated privileges by leveraging the CAP NET ADMIN capability during scan operations with an excessively long SSID value. In this scenario, the lack of proper validation of the SSID length leads to buffer overflows, enabling attackers to exploit the vulnerability, escalate privileges, and execute malicious code. The commit message described that this bug has existed since version 2.6.29-rc4 of the Linux kernel. Given this information, CodeLlama reassesses the risk level of this vulnerability as “High”. This is because it allows attackers to escalate privileges and execute malicious code, and it has persisted for a considerable period of time. It is crucial to address and patch this vulnerability promptly by updating the operating system or kernel to ensure security.
@@
To compare the vulnerability assessment capabilities of LLMs after providing key information, we have created a performance comparison bar chart, as shown in Fig. 6. LLMs have limited capacity for assessing vulnerability severity based solely on source code. However, when provided with key important information, most LLMs (i.e., DeepSeek-Coder, CodeLlama, WizardCoder, and Mistral) exhibit significantly improved vulnerability assessment capabilities, particularly in terms of the Accuracy metric. The Accuracy has increased from the range of 0.26-0.42 to the range of 0.27-0.56. StarCoder and Phi-2 are showing a declining trend, and we believe this may be attributed to the addition of key information, resulting in an increase in the number of input tokens. These LLMs may not excel in handling excessively long text sequences, and we analyze this further in Section 5.2. In contrast, DeepSeek-Coder and Mistral exhibit significant improvements, possibly due to their proficiency in handling long sequential text.
@@
Finding-5. LLMs have the capacity for assessment of vulnerability severity based on source code, and can be improved by providing more context information.
Figure 6: The impact of key important information on LLM Vulnerability Assessment (RQ2)